# SceneSplat: Gaussian Splatting-based Scene Understanding with Vision-Language Pretraining

### v1

@@ 4.3 Self Supervised Pretraining
The proposed GaussianSSL method is presented in Fig. 2 (right). It incorporates multiple losses with different objective into SceneSplat’s large-scale pretraining.
@@
Masked Gaussian Modeling. This part supervises the model to predict masked Gaussian primitives. Given a 3D Gaussian splatting scene represented by $\mathcal{G}=\{G_{i}\}_{i=1}^{N}$, where $G_{i}\in\mathbb{R}^{59}$, the training proceeds as follows. (1) A subset $\{G_{j}\}_{j=1}^{N^{\prime}}$ is sampled from $\mathcal{G}$ using a dense grid sampling of size $S$. (2) Samples are projected into a latent space with the embedding function $P$, in order to obtain tokens $E=P\left(\{G_{j}\}_{j=1}^{N^{\prime}}\right)\in\mathbb{R}^{N^{\prime}\times d_{\text{e}}}$, where $d_{\text{e}}$ is the embedding dimension. (3) The tokens $E$ are masked with ratio $r\in[0,1]$, by replacing $N^{\prime}\cdot r$ randomly chosen tokens with a learnable mask token $t\in\mathbb{R}^{d_{\text{e}}}$, to obtain $T_{\text{m}}\in\mathbb{R}^{N^{\prime}\times d_{\text{e}}}$. (4) The masked tokens $T_{\text{m}}$ are processed with the 3D backbone $g_{\theta}(\cdot)$ to obtain $\hat{T_{\text{m}}}=h_{\phi}(f_{\varphi}(T_{m}))$, where $f_{\varphi}(\cdot)$ is the encoder and $h_{\varphi}(\cdot)$ is the decoder. (5) The output tokens $\hat{T_{\text{m}}}$ are mapped to the input Gaussian space, with the reconstruction projector $\hat{G}_{m}={\Phi}(\hat{T}_{m})\in\mathbb{R}^{N^{\prime}\times F}$. (6) Finally, the $\mathcal{L}_{2}$ reconstruction loss is used between the predicted masked Gaussians and the original ones: $\mathcal{L}_{\text{MGM}}=\mathbb{E}_{G_{j}\sim\mathcal{G}}\left[\|G_{m}-\hat{G}_{m}\|_{2}^{2}\right]$.
@@
Self-Distillation Representation Learning. Self-Distillation [31] learns augmentation-invariant representations by aligning predictions of a student network $\theta_{s}$ with an EMA-updated teacher $\theta_{t}$ [3]. For a batch of Gaussian scenes $\{\mathcal{G}_{n}\}_{n=1}^{B}$ (global/local views $G_{g}^{b}$, $G_{l}^{b}$), we extract tokenized bottleneck features $z\in\mathbb{R}^{M\times d_{e}}$, compute global representations $\bar{z}$ via mean pooling, and align student-teacher outputs via cosine similarity loss $\mathcal{L}_{\text{sim}}$ [52]. Feature diversity is regularized with a coding rate term $\mathcal{L}_{\text{cr}}$ [52, 23] resulting in $\mathcal{L}_{\text{DINO}}=\omega_{\text{sim}}\mathcal{L}_{\text{sim}}+\omega_{\text{cr}}\mathcal{L}_{\text{cr}}$ with corresponding weights. Inspired by [68, 52], the student network also predicts masked feature aligned with the teacher via an $\mathcal{L}_{\text{iBOT}}$, computed using cosine similarity. We propose to mitigate the decoder collapse issues by multi-task reconstruction $\mathcal{L}_{\text{MGM}}$, as coding rate regularization stabilizes only the hierarchical encoder. Following [11, 53], reconstruction is limited to weakly augmented views to avoid degradation.
@@
Language-Gaussian Alignment. As shown in Sec. 4.2, the precomputed language feature enables effective knowledge distillation. For scenes with existing language labels, we seek to leverage them to further regulate self-supervised learning. However, the high dimensionality of these language features (dimension $N\times d_{L}$ ) can substantially increase the computational cost for supervision. To address this, we replace the original features with a compressed representation learned via an autoencoder[19], drastically reducing memory overhead while preserving semantic information. Similar to $\mathcal{L}_{\text{MGM}}$, we use $\mathcal{L}_{\mathrm{LA}}$ follow Eqs. 2 and 3 to train the network to predict low-dimensional language feature from unmasked neighbors.
@@ 5 Experiments
In this section, we evaluate the performance of vision-language pretraining on open-vocabulary task and examine the impact of large-scale Gaussian self-supervised pretraining on downstream indoor semantic segmentation. We further justify our design choices through ablation studies. Implementation details are provided in the supplement.
@@
PLA [8] SN - 1.8 3.1 - - - -
RegionPLC [54] SN - 9.2 16.4 6.2 13.3 11.3 20.1
OV3D [15] SN - 8.7 - - - - -
Mosaic3D [21] SN - 13.0 24.5 8.6 17.8 16.2 27.1
SceneSplat (Ours) SN - 18.9 31.7 10.8 18.7 14.7 24.7
@@
SceneSplat (Ours) SN, SN++ $\times$1.75 19.3 41.5 12.6 23.9 29.1 52.4
@@ 5.1 Vision-Language Pretraining
Tab. 2 reports the zero-shot 3D semantic segmentation results on the fine-grained ScanNet++ (100 classes) [57], Matterport3D (160 classes) [2] and ScanNet200 (200 classes) [5] benchmarks, where methods are trained on specified data sources. When trained on ScanNet, SceneSplat achieves the state-of-the-art results, leading to a $5.9\%$ and $2.2\%$ f-mIoU increase on the ScanNet200 and Matterport3D benchmarks respectively. By extending the training sources, we obtain a $3.6\%$ and $11.1\%$ f-mIoU increase on the ScanNet200 and ScanNet++ benchmark, falling behind only by $0.5\%$ f-mIoU compared to concurrent work on Matterport3D. Notably, [21] uses 14 times more training scenes to achieve its best results.
Fig. 3 shows the zero-shot segmentation results on evaluation scenes. SceneSplat not only achieves competitive segmentation performance but also correctly annotates the missing semantic labels (e.g., desks shown above). We demonstrate text-based queries on the inference results of predicted language features in Fig. 4. Our vision-language pretraining enables effective localization of complex objects within the scene.
@@ 5.2 Label-free 3DGS Pretraining
We conduct extensive pretraining experiments with the proposed GaussianSSL method. To assess the efficacy of the pretrained model, we report the segmentation results in Tab. 3. Our method achieves a +0.1% improvement over supervised-only baselines on ScanNet20 and +0.5% on ScanNet200, while observing a performance drop on ScanNet++ primarily due to pretraining dataset quality variations (Tab. 1). Furthermore, compared to our reproduced implementation of PTv3  [51], we outperform by +0.8% on ScanNet20 and +0.9% on ScanNet200 (Tab. 4). More qualitative results are in the supplement.
@@ 5.3 Further Statistical Evaluation
SceneSplat Inference Results vs. Collected Language Labels. One may assume the collected language labels cap the upper bound zero-shot performance since they provide the supervision signal during vision-language pretraining. Interestingly, this is not always the case. Tab. 5 compares the performance of SceneSplat inference features with the collected language labels. On ScanNet++, ours outperforms the labels with a $4.2\%$ increase in f-mIoU. Although the collected labels are not perfect, large-scale pretraining is able to filter noise and learn meaningful patterns.
@@
Impact of the Input 3DGS Scene PSNR on Open-Vocabulary Performance. Reported on the Matterport3D test split with 370 scenes, Fig. 5 indicates a clear positive trend of the input 3DGS scene PSNR of the training views and the resulted zero-shot mIoU performance with SceneSplat model. Low PSNRs usually come out of blurry input images, poor Gaussian centers optimization and insufficient scene coverage, where the 3DGS parameters cannot well resolve the scene. The trend here highlights the importance of data curation on the collected 3DGS scenes.
@@
Nearest Neighbors Voting During Zero-Shot Experiments. The centers of input Gaussians differ from the point locations where the semantic predictions are evaluated on, thus we have to aggregate predictions from neighboring Gaussians. We carry out majority voting using the nearest neighboring Gaussians for each evaluating location. Fig. 6 ablates the number of nearest neighbors on the IoU results using the ScanNet++ validation split. We observe the overall trend of mIoU increase w.r.t. to the number of nearest neighbors and list the classes with top relative changes. To balance the performance and inference speed, 25 nearest neighbors are selected for voting.
@@
Ablation on Contrastive Loss in the Lang-Aware Pretraining. Tab. 7 ablates the contrastive loss applied during vision-language pretraining, where applying contrastive loss at the late stage of training outperforms other variants.
Runtime vs. Per-Scene Language Gaussian Splatting Method. Thanks to the feed-forward ability after vision-language pretraining, SceneSplat is shown in Tab. 8 to be 445.8 times faster than the state-of-the-art Language-embedded Gaussian Splatting method, as there is no need for 2D feature extraction and fusion.
@@ 6 Conclusion
In this work, we introduce SceneSplat, the first large-scale 3D scene understanding model for indoor environments operating directly on 3D Gaussian splats. Powered by SceneSplat-7K, a dataset comprising 6,868 scenes, we propose a novel 3D Gaussian splat encoder that generates semantic features in a single pass, enabling open-vocabulary scene recognition without relying on 2D fusion. Through self-supervised techniques, we unlock label-free 3DGS pretraining at the scene level. Our approach achieves state-of-the-art performance in zero-shot semantic segmentation, establishing new benchmarks and laying the foundation for future advancements in open-vocabulary 3D understanding.
@@ A.1 Language Label Collection
Many existing methods align 3D primitives with text embeddings from vision-language models (VLM) [8, 21]. While effective for basic understanding, this approach inherently limits the information captured, as textual descriptions typically only convey categorical and spatial properties, lacking fine-grained visual details and relationships. Methods employing visual captioning models [26, 59] face similar challenges, as they struggle to describe all aspects of the target scene content. Even detailed captions inevitably result in information loss during the text embedding alignment process. Differently, we directly align Gaussians with the image embedding space of VLM, thereby preserving richer latent semantic information.
Dynamic Weighting Mechanism. Unlike previous approaches that use a single tight crop around each segment, we adapt the three-crop strategy [49] with dynamic weighting to capture the context. For each segment identified by SAMv2, we extract three distinct features: (1) Global feature $f_{g}$ from the entire RGB frame, capturing the full scene context; (2) Local feature $f_{l}$ from the image crop with background; (3) Masked feature $f_{m}$ from the image crop without background, focusing solely on the object. We use SAMv2/sam2-hiera-large and SigLIP2/siglip2-base-patch16-512 models.
@@
The final fused feature is computed as a weighted combination and normalized:
@@
This mechanism adaptively balances global context influence via $w_{g}$, local context with background via $w_{l}$, and object-specific features via $w_{m}$. These three features are dynamically combined to create a representation that adapts based on the segment’s relationship with its context. For objects highly integrated with their surroundings (e.g., a keyboard in front of a monitor), background-inclusive features receive a higher weight. For isolated objects (e.g., a coffee mug), background-excluded features dominate.
@@ A.2 Vision-Language Pretraining
Our vision-language 3DGS pretraining model is built on a transformer encoder-decoder architecture adapted from [51]. As detailed in Tab. F, the encoder consists of 4 stages with depths [2,2,2,6], channels [32,64,128,256], and heads [2,4,8,16], while the decoder employs 3 stages with depths [2,2,2], channels [768,512,256], and 16 attention heads each. We process 3D Gaussian primitives with all parameters (center, color, opacity, quaternion, and scale). The model is trained with AdamW optimizer (initial LR = 0.006, weight decay = 0.05) using a OneCycle scheduler with cosine annealing. Our loss weights are set to $\lambda_{\text{cos}}=1.0$, $\lambda_{\text{L2}}=1.0$, and $\lambda_{\text{con}}=0.02$, with contrastive loss (temperature $\tau=0.2$) activated only in the later $75\%$ of training. During training, we employ extensive data augmentation including random rotations, scaling, flipping, jittering, and elastic distortions (see Tab. H). We train for 800 data epochs using 4 NVIDIA H100 GPUs.
@@ A.3 Masked Gaussian Modeling
For all self-supervised pretraining experiments on the full dataset, we employ 4 NVIDIA H100 GPUs (80GB each) and the AdamW optimizer, with learning rates of $1\times 10^{-3}$ for the embedding layer and $1\times 10^{-4}$ for the attention blocks, coupled with a weight decay of $1\times 10^{-3}$. We use mixed-precision training (FP16) to accelerate convergence and reduce memory overhead. In total we train for 300k steps. We use an $\mathcal{L}_{2}$ loss to enforce consistency between predicted and ground truth Gaussian parameters, focusing only on masked regions. For parameter reconstruction, we design a three-layer MLP projector for each Gaussian attribute, incorporating output activations that adhere to physical constraints: Color use a $\tanh$ activation (bounded in $[-1,1]$), Opacity and Scale use a sigmoid activation(bounded in $[0,1]$) because in most indoor scenes, most Gaussians have small scales, Rotation, and Normals applies $\tanh$ followed by $\ell_{2}$-normalization to ensure valid quaternions and normals.

### v2

@@ 4.3 Self Supervised Pretraining
The proposed GaussianSSL method is presented in Fig. 2 (right). It incorporates multiple losses with different objectives into SceneSplat’s large-scale pretraining.
@@
Masked Gaussian Modeling. This part supervises the model to predict masked Gaussian primitives. Given a 3D Gaussian splatting scene represented by $\mathcal{G}=\{G_{i}\}_{i=1}^{N}$, where $G_{i}\in\mathbb{R}^{59}$, the training proceeds as follows: (1) A subset $\{G_{j}\}_{j=1}^{N^{\prime}}$ is sampled from $\mathcal{G}$ using dense grid sampling of size $S$. (2) Samples are projected into a latent space with the embedding function $P$ to obtain tokens $E=P\left(\{G_{j}\}_{j=1}^{N^{\prime}}\right)\in\mathbb{R}^{N^{\prime}\times d_{\text{e}}}$, where $d_{\text{e}}$ is the embedding dimension. (3) The tokens $E$ are masked with ratio $r\in[0,1]$ by replacing $N^{\prime}\cdot r$ randomly chosen tokens with a learnable mask token $t\in\mathbb{R}^{d_{\text{e}}}$ to obtain $T_{\text{m}}\in\mathbb{R}^{N^{\prime}\times d_{\text{e}}}$. (4) The masked tokens $T_{\text{m}}$ are processed using the 3D backbone $g_{\theta}(\cdot)$ to obtain $\hat{T_{\text{m}}}=h_{\phi}(f_{\varphi}(T_{m}))$, where $f_{\varphi}(\cdot)$ is the encoder and $h_{\varphi}(\cdot)$ is the decoder. (5) The output tokens $\hat{T_{\text{m}}}$ are mapped to the input Gaussian space with the reconstruction projector $\hat{G}_{m}={\Phi}(\hat{T}_{m})\in\mathbb{R}^{N^{\prime}\times F}$. (6) Finally, the $\mathcal{L}_{2}$ reconstruction loss between the predicted masked Gaussians and the original Gaussians is used: $\mathcal{L}_{\text{MGM}}=\mathbb{E}_{G_{j}\sim\mathcal{G}}\left[\|G_{m}-\hat{G}_{m}\|_{2}^{2}\right]$.
@@
Self-Distillation Representation Learning. Self-distillation [31] learns augmentation-invariant representations by aligning the predictions of a student network $\theta_{s}$ with an EMA-updated teacher $\theta_{t}$ [3]. For a batch of Gaussian scenes $\{\mathcal{G}_{n}\}_{n=1}^{B}$ (global/local views $G_{g}^{b}$, $G_{l}^{b}$), we extract tokenized bottleneck features $z\in\mathbb{R}^{M\times d_{e}}$, compute global representations $\bar{z}$ via mean pooling, and align student-teacher outputs via cosine similarity loss $\mathcal{L}_{\text{sim}}$ [52]. Feature diversity is regularized with a coding rate term $\mathcal{L}_{\text{cr}}$ [52, 23] resulting in $\mathcal{L}_{\text{DINO}}=\omega_{\text{sim}}\mathcal{L}_{\text{sim}}+\omega_{\text{cr}}\mathcal{L}_{\text{cr}}$ with corresponding weights. Inspired by [68, 52], the student network also predicts masked features aligned with the teacher via $\mathcal{L}_{\text{iBOT}}$, computed using cosine similarity. We propose to mitigate the decoder collapse issues by multi-task reconstruction $\mathcal{L}_{\text{MGM}}$, as coding rate regularization stabilizes only the hierarchical encoder. Following [11, 53], the reconstruction is limited to weakly augmented views to avoid degradation.
@@
Language-Gaussian Alignment. As shown in Sec. 4.2, the precomputed language feature enables effective knowledge distillation. For scenes with existing language labels, we seek to leverage them to further regulate self-supervised learning. However, the high dimensionality of these language features (dimension $N\times d_{L}$ ) can substantially increase the computational cost of supervision. To address this, we replace the original features with a compressed representation learned via an autoencoder [19], drastically reducing the memory overhead while preserving semantic information. Similar to $\mathcal{L}_{\text{MGM}}$, we use $\mathcal{L}_{\mathrm{LA}}$ following Eqs. 2 and 3 to train the network to predict low-dimensional language features from unmasked neighbors.
@@ 5 Experiments
In this section, we evaluate the performance of vision-language pretraining on open-vocabulary task and examine the impact of large-scale Gaussian self-supervised pretraining on downstream indoor semantic segmentation. We further justify our design choices through ablation studies. The implementation details are provided in the supplement.
@@
PLA [8] SN 1.8 3.1
RegionPLC [54] SN 9.2 16.4 6.2 13.3 11.3 20.1
OV3D [15] SN 8.7
Mosaic3D [21] SN 13.0 24.5 8.6 17.8 16.2 27.1
SceneSplat (Ours) SN 18.9 31.7 10.8 18.7 14.7 24.7
@@
SceneSplat (Ours) SN, SN++, MP3D $\times$2.92 21.4 38.7 13.8 31.8 28.4 50.0
@@ 5.1 Vision-Language Pretraining
Tab. 2 reports the zero-shot 3D semantic segmentation results on the fine-grained ScanNet++ (100 classes) [57], Matterport3D (160 classes) [2] and ScanNet200 (200 classes) [5] benchmarks, where methods are trained on specified data sources. When trained on ScanNet, SceneSplat achieves state-of-the-art results, leading to $5.9\%$ and $2.2\%$ f-mIoU increases on the ScanNet200 and Matterport3D benchmarks. By extending the training sources, we obtain $5.7\%$, $0.7\%$, and $10.4\%$ f-mIoU increases on the ScanNet200, Matterport3D, and ScanNet++ benchmarks, respectively, compared to concurrent work [21]. Notably, [21] uses 8.32$\times$ training scenes to achieve its best results.
Fig. 3 shows the zero-shot segmentation results on evaluation scenes. SceneSplat not only achieves competitive segmentation performance but also correctly annotates the missing semantic labels (e.g., desks shown above). We demonstrate text-based queries on the inference results of the predicted language features in Fig. 4. Our vision-language pretraining enables the effective localization of complex objects within the scene.
@@ 5.2 Label-free 3DGS Pretraining
We conduct extensive pretraining experiments with the proposed GaussianSSL method. To assess the efficacy of the pretrained model, we report the segmentation results in Tab. 3. Our method achieves a +0.1% improvement over supervised-only baselines on ScanNet20 and +0.5% on ScanNet200, while observing a performance drop on ScanNet++ primarily due to pretraining dataset quality variations (Tab. 1). Furthermore, compared with our reproduced implementation of PTv3  [51], we outperform by +0.8% on ScanNet20 and +0.9% on ScanNet200 (Tab. 4). More qualitative results are provided in the supplement.
@@ 5.3 Further Statistical Evaluation
SceneSplat Inference Results vs. Collected Language Labels. One may assume that the collected language labels cap the upper bound zero-shot performance since they provide the supervision signal during vision-language pretraining. Interestingly, this is not always the case. Tab. 5 compares the performance of SceneSplat inference features with the collected language labels. On ScanNet++, our method outperforms the labels with a $4.2\%$ increase in f-mIoU. Although the collected labels are not perfect, large-scale pretraining can filter noise and learn meaningful patterns.
@@
Impact of the Input 3DGS Scene PSNR on Open-Vocabulary Performance. Reported on the Matterport3D test split with 370 scenes, Fig. 5 indicates a clear positive trend of the input 3DGS scene PSNR of the training views and the resulting zero-shot mIoU performance with SceneSplat model. Low PSNRs usually come out of blurry input images, poor Gaussian centers optimization, and insufficient scene coverage, where the 3DGS parameters cannot resolve the scene well. This trend highlights the importance of data curation for the collected 3DGS scenes.
@@
Nearest Neighbors Voting During Zero-Shot Experiments. The centers of the input Gaussians differ from the point locations where the semantic predictions are evaluated; thus, we have to aggregate predictions from neighboring Gaussians. We perform majority voting using the nearest neighboring Gaussians for each evaluation location. Fig. 6 ablates the number of nearest neighbors on the IoU results using the ScanNet++ validation split. We observe the overall trend of the mIoU increase w.r.t. to the number of nearest neighbors and list the classes with the top relative changes. To balance the performance and inference speed, 25 nearest neighbors are selected for voting.
@@
Ablation on Contrastive Loss in the Vision-Language Pretraining. Tab. 7 ablates the contrastive loss applied during vision-language pretraining, where applying contrastive loss at the late stage of training outperforms other variants.
Runtime vs. Per-Scene Language Gaussian Splatting Method. Thanks to the feed-forward ability after vision-language pretraining, SceneSplat is shown in Tab. 8 to be 445.8$\times$ faster than the state-of-the-art language-embedded Gaussian Splatting method, as there is no need for 2D feature extraction and fusion.
@@ 6 Conclusion
In this work, we introduce SceneSplat, the first large-scale 3D scene understanding model for indoor environments operating directly on 3D Gaussian splats. Powered by SceneSplat-7K, a dataset comprising 7,916 scenes, we propose a novel 3D Gaussian splat encoder that generates semantic features in a single pass, enabling open-vocabulary scene recognition without relying on 2D fusion. Through self-supervised techniques, we unlock label-free 3DGS pretraining at the scene level. Our approach achieves state-of-the-art performance in zero-shot semantic segmentation, establishing new benchmarks and laying the foundation for future advancements in open-vocabulary 3D understanding.
7 Acknowledgment
Yue Li is financially supported by TomTom, the University of Amsterdam and the allowance of Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy. This work used Dutch national e-infrastructure with the support of the SURF Cooperative under grant no. NWO-2024.035. This work was partially supported by INSAIT, Sofia University “St. Kliment Ohridski” (Partially funded by the Ministry of Education and Science of Bulgaria’s support for INSAIT as part of the Bulgarian National Roadmap for Research Infrastructure), the MUR PNRR project FAIR (PE00000013), the EU Horizon project ELIAS (No. 101120237), and the computational resources provided by the Google Cloud Platform (GCP).
@@ A.1 Language Label Collection
Many existing methods align 3D primitives with text embeddings from vision-language models (VLM) [8, 21]. While effective for basic understanding, this approach inherently limits the information captured, as textual descriptions typically only convey categorical and spatial properties, lacking fine-grained visual details and relationships. Methods employing visual captioning models [26, 59] face similar challenges as they struggle to describe all aspects of the target scene content. Even detailed captions inevitably result in information loss during the text embedding alignment process. In contrast, we directly align Gaussians with the image embedding space of VLM, thereby preserving richer latent semantic information.
Dynamic Weighting Mechanism. Unlike previous approaches that use a single tight crop around each segment, we adapt the three-crop strategy [49] with dynamic weighting to capture the context. For each segment identified by SAMv2, we extract three distinct features: (1) global feature $f_{g}$ from the entire RGB frame, capturing the full scene context; (2) local feature $f_{l}$ from the image crop with background; (3) masked feature $f_{m}$ from the image crop without background, focusing solely on the object. During the process, SAMv2/sam2-hiera-large and SigLIP2/siglip2-base-patch16-512 models are used.
@@
The final fused feature is computed as a weighted combination and normalized as:
@@
This mechanism adaptively balances the global context influence via $w_{g}$, local context with background via $w_{l}$, and object-specific features via $w_{m}$. These three features are dynamically combined to create a representation that adapts to the segment’s relationship with its context. For objects that are highly integrated with their surroundings (e.g., a keyboard in front of a monitor), background-inclusive features receive a higher weight. For isolated objects (e.g., a coffee mug), background-excluded features dominate.
@@ A.2 Vision-Language Pretraining
Our vision-language 3DGS pretraining model is built on a transformer encoder-decoder architecture adapted from [51]. As detailed in Tab. A, the encoder consists of 4 stages with depths [2,2,2,6], channels [32,64,128,256], and heads [2,4,8,16], while the decoder employs 3 stages with depths [2,2,2], channels [768,512,256], and 16 attention heads each. We process 3D Gaussian primitives with all parameters (center, color, opacity, quaternion, and scale). The model is trained with the AdamW optimizer (initial LR = 0.006, weight decay = 0.05) using a OneCycle scheduler with cosine annealing. Our loss weights are set to $\lambda_{\text{cos}}=1.0$, $\lambda_{\text{L2}}=1.0$, and $\lambda_{\text{con}}=0.02$, with contrastive loss (temperature $\tau=0.2$) activated only in the later $75\%$ of training. During training, we employ extensive data augmentation, including random rotations, scaling, flipping, jittering, and elastic distortions (see Tab. C). We train for 800 data epochs using 4 NVIDIA H100 GPUs.
@@ A.3 Masked Gaussian Modeling
For all self-supervised pretraining experiments on the full dataset, we employ 4 NVIDIA H100 GPUs (94GB each) and the AdamW optimizer, with learning rates of $1\times 10^{-3}$ for the embedding layer and $1\times 10^{-4}$ for the attention blocks, coupled with a weight decay of $1\times 10^{-3}$. We use mixed-precision training (FP16) to accelerate convergence and reduce memory overhead. In total, we train for 300k steps. We use an $\mathcal{L}_{2}$ loss to enforce consistency between predicted and ground truth Gaussian parameters, focusing only on masked regions. For parameter reconstruction, we design a three-layer MLP projector for each Gaussian attribute, incorporating output activations that adhere to physical constraints: color uses a $\tanh$ activation (bounded in $[-1,1]$), opacity and scale use a sigmoid activation (bounded in $[0,1]$) because in most indoor scenes, most Gaussians have small scales, rotation, and normals apply $\tanh$ followed by $\ell_{2}$-normalization to ensure valid quaternions and normals.